# Speak, Segment, Track, Navigate: An Interactive System for Video-Guided Skull-Base Surgery

### 17 Mar 2026 v1

@@ Abstract
We introduce a speech-guided embodied agent framework for video-guided skull base surgery that dynamically executes perception and image-guidance tasks in response to surgeon queries. The proposed system integrates natural language interaction with real-time visual perception directly on live intraoperative video streams, thereby enabling surgeons to request computational assistance without disengaging from operative tasks. Unlike conventional image-guided navigation systems that rely on external optical trackers and additional hardware setup, the framework operates purely on intraoperative video. The system begins with interactive segmentation and labeling of the surgical instrument. The segmented instrument is then used as a spatial anchor that is autonomously tracked in the video stream to support downstream workflows, including anatomical segmentation, interactive registration of preoperative 3D models, monocular video-based estimation of the surgical tool pose, and support image guidance through real-time anatomical overlays. We evaluate the proposed system in video-guided skull base surgery scenarios and benchmark its tracking performance against a commercially available optical tracking system. Across three experimental trials, the vision-based method achieved a mean absolute tool-tip position error of $2.83\pm 1.64$ mm and roll and pitch errors of $0.16\pm 0.21^{\circ}$ and $0.17\pm 0.22^{\circ}$, respectively. The system completes tool segmentation and anatomy registration within approximately two minutes, substantially reducing setup complexity compared to conventional tracking workflows. These results demonstrate that speech-guided embodied agents can achieve competitive spatial accuracy while improving workflow integration and enabling rapid deployment of video-guided surgical systems.
@@ I Introduction
Interactive segmentation frameworks based on vision foundation models (VFMs) partially address usability through prompt-driven paradigms. Models such as the SAM [7] enable category-agnostic segmentation via visual prompts, while Grounded SAM [19] extends this paradigm to text-conditioned segmentation. Although these approaches represent a shift toward more flexible human-machine interaction, they typically require direct visual inputs (e.g., clicks, strokes, or bounding boxes) or static textual prompts. They are not inherently designed for continuous, hands-free operation within dynamic surgical environments, nor do they orchestrate multiple downstream tasks beyond segmentation. Recent advances in large language models (LLMs) and vision–language models (VLMs) have catalyzed the emergence of AI agents for surgical applications. Systems such as TPSIS [29] and RSVIS [23] incorporate multimodal reasoning for semantically guided segmentation, emulate collaborative clinical roles, SurgRAW [13] coordinates multiple vision–language agents using structured reasoning, LLaVA-Surge [10] addresses open-ended video understanding, and SuFIA [16] interprets surgical scenes while planning context-aware responses. These approaches rely heavily on the visual perception capabilities of VLMs. While expressive, VLMs have two major limitations: 1) these models require extensive finetuning and vocabulary grounding for new tasks or surgical domains. 2) they struggle to perform online visual perception tasks in dynamic realworld environments [4]. These models contain billions of parameters, thereby creating computational and memory requirements that exceed the capabilities of edge devices where embodied agents typically operate.
@@
In this work, we propose SCOPE, a speech-guided embodied agent for enabling effective surgical assistance in video-guided skull base surgery. The proposed system is capable of interacting naturally with surgeons and execute perception and image-guidance tasks directly on live intraoperative video streams, thereby enabling hands-free operation without disengaging from the operative field. The system begins with interactive segmentation and labeling of the surgical instrument, which subsequently serves as a spatial anchor for downstream operations. By autonomously tracking the instrument within the video stream, the agent enables additional workflows including anatomical segmentation, registration of preoperative 3D models, and monocular video-based estimation of the surgical tool pose. These capabilities are coordinated through a conversational interface that allows surgeons to dynamically request computational assistance during the procedure. Rather than relying on monolithic multimodal VLM models, we adopt a two-stage approach that decouples reasoning and perception, where LLM can focus on intent understanding and task orchestration, while specialized vision foundation models perform spatially precise perception tasks such as segmentation and tracking. This separation enables a modular architecture in which components can be independently upgraded, reduces computational overhead, and facilitates deployment on edge hardware equipped with commercial-grade GPUs. Furthermore, instead of relying on task-specific fine-tuning with surgery-specific datasets, which are often limited in size and diversity, we prioritize the use of zero-shot VFMs. Combined with human-in-the-loop interaction, this strategy allows rapid adaptation to new surgical contexts while avoiding overfitting and costly retraining.
@@
We evaluate the proposed system in video-guided skull base surgery scenarios and benchmark its 3D tracking performance against a commercially available optical tracking system. Across three experimental trials, the vision-based method achieved a mean absolute tool-tip position error of $2.83\pm 1.64$ mm and roll and pitch errors of $0.16\pm 0.21^{\circ}$ and $0.17\pm 0.22^{\circ}$, respectively. The system can complete tool segmentation and anatomy registration within approximately two minutes, which substantially reduces setup complexity compared to conventional tracking workflows. These results demonstrate that a modular, speech-guided embodied agent can achieve competitive spatial accuracy while enabling rapid deployment and improved workflow integration for video-guided surgical navigation.
@@ II Related Work
Tool pose estimation under symmetry and limited observability: Recent surgical-tool pose estimation methods increasingly rely on learning-based pipelines, including task-specific monocular pose networks and joint detection–segmentation–geometry formulations tailored to surgical scenes [9, 5]. In parallel, the broader 6D pose literature has developed model-based correspondence methods, where dense or sparse image-to-model correspondences are predicted and then solved with PnP, as in ZebraPose [21], as well as model-free approaches such as Any6D that relax the requirement for a fixed object model. While these approaches are powerful, surgical video remains especially challenging because elongated instruments often exhibit weak texture, specular highlights, partial occlusion, foreshortening, and near-symmetry, all of which reduce pose observability. Prior work in robotic surgery has therefore also explored incorporating additional priors, such as robot kinematics and temporal filtering, to improve robustness [17].
Our pose tracking formulation is designed for this regime. Rather than relying solely on appearance-based correspondence prediction, we combine (i) mask-derived 2D axis constraints, (ii) depth-based 3D priors, and (iii) hypothesis gating via silhouette agreement to resolve ambiguities for elongated, near-symmetric tools. In addition, we use a foreground-conditioned monocular depth pipeline and recover metric scale using registered anatomy depth, thereby coupling learned depth with geometric consistency for stable reconstruction and tracking. Overall, while prior work provides strong building blocks for multimodal assistance and promptable perception, our contribution is a unified real-time embodied workflow that couples speech-driven control with continuous segmentation, temporally stable tip tracking, trajectory-conditioned anatomy prompting, interactive PnP registration, and depth-aware navigation/pose tracking on live surgical video.
Figure 1: System overview of the embodied surgical agent. The surgeon interacts through a hands-free interface (speech-to-text) that issues high-level commands to the front end, which orchestrates live video streaming, tool/anatomy segmentation, pose tracking, and anatomy registration. Intermediate outputs (masks, pose hypotheses, and registered anatomy overlays) are persisted in a streaming memory and can be retrieved on demand to support iterative refinement and rapid task switching without disrupting the surgical workflow. The back end executes modular perception and geometry components—including promptable segmentation, temporal mask propagation, surgical navigation, and pose/registration solvers—to produce stable tool state and anatomy-aligned navigation overlays in the endoscopic view.
@@ III Methodology
III-A Surgical Tool Segmentation
We build on the speech-guided tool segmentation pipeline introduced in the previous work  [14], which combines open-vocabulary detection and promptable segmentation (GroundingDINO+SAM / GSAM) for mask proposal generation and CUTIE for temporal propagation [14, 12, 7, 19, 3]. This approach’s major limitation is that the tool must remain approximately stationary during human selection of the desired mask because propagation begins only after tracking is enabled.
To remove this interruption, we introduce a streaming memory module that buffers incoming frames together with previously segmented entities while the surgeon interacts. Once a tool mask is confirmed at time $t_{0}$, the selected mask is stored and used to seed a catch-up propagation step over a small set of uniformly sampled buffered frames up to the current time $t_{n}$. This synchronizes the tracker with the live scene and allows real-time mask propagation to resume without requiring the tool to pause during selection.
@@ III-A Surgical Tool Segmentation
III-B Tip Point Tracking
Previous work [14] tracked the surgical tool tip using the principal axis of the tool mask obtained by two-dimensional principal component analysis (2D PCA). While effective under full visibility, this formulation is sign-ambiguous and becomes unstable when the tool is truncated or partially occluded.
We therefore enforce temporal consistency on the PCA endpoints. Let $m_{t}$ denote the binary tool mask at time $t$, and let $\mathbf{e}_{t}^{(1)}$ and $\mathbf{e}_{t}^{(2)}$ denote the two extrema of the mask projection along its first principal axis. At initialization, the endpoint closer to the image boundary is treated as the tool base and the opposite endpoint as the tip. For subsequent frames, the tip is selected as the endpoint closest to the previous estimate:
@@ III-B Tip Point Tracking
This suppresses axis flips and preserves directional consistency over time.
To reduce PCA bias under partial visibility, we further apply a base-aware boundary-cropping step. If the estimated base lies on or near the image boundary, we iteratively remove a small strip of the mask on the base side and recompute PCA until the base is no longer boundary-truncated. This yields a more stable principal-axis estimate under partial views.
III-C Anatomy Segmentation
Figure 3: Interactive anatomy segmentation and refinement with event-triggered prompt retrieval. A voice command starts anatomy segmentation while the system buffers the drill-tip trajectory. When the surgeon says “Done,” the stored trajectory is converted into spatial prompts for SAM to generate an initial mask. Refinement is triggered by a “Click” command over regions to remove, adding prompts for mask update. The final anatomy mask $M^{\mathrm{Anat}}$ seeds CUTIE for temporal propagation. Blue arrows denote segmentation-mode prompting and trajectory retrieval; purple arrows denote refinement interactions and SAM updates.
While surgical tools can be segmented effectively using vision–language models, anatomy segmentation remains more challenging because of non-rigid geometry and appearance variability across surgical phases. Previous work [14] addressed the impracticality of cursor-based interaction by introducing hands-free anatomy segmentation with a virtual cursor anchored to the instrument tip. In that framework, monocular depth from DepthAnything [26] is used to trigger a virtual click near the tip, and the resulting point prompt is passed to SAM [7] to generate an anatomy mask for temporal propagation. However, this point-based interaction is sensitive to monocular depth scale ambiguity and often provides weak spatial constraints, leading to over- or under-segmentation.
To overcome these limitations, we introduce region-constrained anatomy segmentation, which encodes user intent through spatial constraints derived from tool motion. As the user hovers the instrument over a region of interest, we collect a trajectory $\mathcal{T}=\{\mathbf{u}_{i}\}_{i=1}^{N}$, where $\mathbf{u}_{i}\in\mathbb{R}^{2}$ denotes the projected tool tip location in the image plane. A bounding box $B=[x_{\min},x_{\max}]\times[y_{\min},y_{\max}]$ is computed from $\mathcal{T}$, defining a constrained interaction region. We then sample a set of point prompts
@@ III-E Surgical Navigation
III-F Spatial Pose Tracking
Accurate spatial pose tracking of surgical instruments is essential for navigation and anatomy-aware guidance [22, 11]. Although state of the art pose estimators such as FoundationPose [24] are effective for general objects, they are less reliable in surgical video, where instruments are elongated, nearly symmetric, and frequently affected by specularities and partial occlusions. These factors make hypothesis refinement unstable and can lead to inconsistent convergence or axis flips.
We therefore formulate pose estimation as a constrained tracking problem tailored to elongated surgical tools. The 2D segmentation mask provides the image-space principal axis, while depth recovers the 3D axis direction and tip location in the camera frame. Temporal consistency is then enforced through axis continuity constraints, enabling more stable pose estimates under partial visibility and large rotations.
Fore-ground Conditioned Depth Retrieval
Accurate 3D reconstruction of the tool axis and tip requires reliable foreground depth. We found stereo-based methods, including IGEV++ and FoundationStereo, to degrade under high magnification because of specularities, low-texture anatomy, and narrow effective baseline. In contrast, monocular depth provides denser and more stable relative predictions. Empirically, the most reliable metric depth recovery was obtained by combining fore-ground conditioned DepthAnything v2 (DA2) [26] with depth from the registered anatomical model.
Figure 6: Stereo vs. fore-ground conditioned monocular depth for surgical tool reconstruction. Stereo depth estimates degrade under specularities and low texture, leading to noisy 3D reconstructions, whereas DepthAnything v2 constrained to the segmented fore-ground produces denser, more consistent depth and a smoother tool point cloud.
Let frames be indexed by $t\in\{1,\dots,N\}$, and pixels by $\mathbf{u}=(u,v)$. After anatomy registration, we render a metric depth map directly from the registered 3D anatomical mesh under the estimated camera pose:
$S_{f}^{\mathrm{reg}}(\mathbf{u}),$
which provides reliable metric depth over visible anatomical regions.
Let the foreground mask be defined as the union of all propagated object masks:
$M_{f}^{\mathrm{fg}}(\mathbf{u})=M_{f}^{\mathrm{tool}}(\mathbf{u})\;\vee\;M_{f}^{\mathrm{anat}}(\mathbf{u}),$
where $M_{f}^{\mathrm{tool}}$ and $M_{f}^{\mathrm{anat}}$ denote the tool and anatomy masks, respectively.

### 16 Apr 2026 v2

# Speak, Segment, Track, Navigate: An Interactive System for Monocular Video-Guided Skull-Base Surgery

@@ Abstract
We introduce a speech-guided embodied agent framework for video-guided skull base surgery that dynamically executes perception and image-guidance tasks in response to surgeon queries. The proposed system integrates natural language interaction with real-time visual perception directly on live intraoperative video streams, thereby enabling surgeons to request computational assistance without disengaging from operative tasks. Unlike conventional image-guided navigation systems that rely on external optical trackers and additional hardware setup, the framework operates purely on intraoperative video. The system begins with interactive segmentation and labeling of the surgical instrument. The segmented instrument is then used as a spatial anchor that is autonomously tracked in the video stream to support downstream workflows, including anatomical segmentation, interactive registration of preoperative 3D models, monocular video-based estimation of the surgical tool pose, and support image guidance through real-time anatomical overlays. We evaluate the proposed system in video-guided skull base surgery scenarios and benchmark its tracking performance against a commercially available optical tracking system. Across three experimental trials, the hybrid vision-based method achieved a mean absolute tool-tip position error of $2.32\pm 1.10$ mm in the camera frame, with inter-frame yaw and pitch propagation discrepancies of $0.18\pm 0.25^{\circ}$ and $0.21\pm 0.30^{\circ}$, respectively. The system completes tool segmentation and anatomy registration within approximately two minutes, substantially reducing setup complexity relative to conventional tracking workflows. These results demonstrate that speech-guided embodied agents can provide accurate spatial guidance while improving workflow integration and enabling rapid deployment of video-guided surgical systems.
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ I Introduction
Interactive segmentation frameworks based on vision foundation models (VFMs) partially address usability through prompt-driven paradigms. Models such as the SAM [9] enable category-agnostic segmentation via visual prompts, while Grounded SAM [18] extends this paradigm to text-conditioned segmentation. Although these approaches represent a shift toward more flexible human-machine interaction, they typically require direct visual inputs (e.g., clicks, strokes, or bounding boxes) or static textual prompts. They are not inherently designed for continuous, hands-free operation within dynamic surgical environments, nor do they orchestrate multiple downstream tasks beyond segmentation. Recent advances in large language models (LLMs) and vision–language models (VLMs) have catalyzed the emergence of AI agents for surgical applications. Systems such as TPSIS [29] and RSVIS [22] incorporate multimodal reasoning for semantically guided segmentation, emulate collaborative clinical roles, SurgRAW [14] coordinates multiple vision–language agents using structured reasoning, LLaVA-Surge [11] addresses open-ended video understanding, and SuFIA [16] interprets surgical scenes while planning context-aware responses. These approaches rely heavily on the visual perception capabilities of VLMs.
While expressive, VLMs have two major limitations: 1) these models require extensive finetuning and vocabulary grounding for new tasks or surgical domains. 2) they struggle to perform online visual perception tasks in dynamic realworld environments [5]. These models contain billions of parameters, thereby creating computational and memory requirements that exceed the capabilities of edge devices where embodied agents typically operate.
@@
In this work, we propose a speech-guided embodied agent for enabling effective surgical assistance in video-guided skull base surgery. The proposed system is capable of interacting naturally with surgeons and execute perception and image-guidance tasks directly on live intraoperative video streams, thereby enabling hands-free operation without disengaging from the operative field. The system begins with interactive segmentation and labeling of the surgical instrument, which subsequently serves as a spatial anchor for downstream operations. By autonomously tracking the instrument within the video stream, the agent enables additional workflows including anatomical segmentation, registration of preoperative 3D models, and monocular video-based estimation of the surgical tool pose. These capabilities are coordinated through a conversational interface that allows surgeons to dynamically request computational assistance during the procedure. Rather than relying on monolithic multimodal VLM models, we adopt a two-stage approach that decouples reasoning and perception, where LLM can focus on intent understanding and task orchestration, while specialized vision foundation models perform spatially precise perception tasks such as segmentation and tracking. This separation enables a modular architecture in which components can be independently upgraded, reduces computational overhead, and facilitates deployment on edge hardware equipped with commercial-grade GPUs. Furthermore, instead of relying on task-specific fine-tuning with surgery-specific datasets, which are often limited in size and diversity, we prioritize the use of zero-shot VFMs. Combined with human-in-the-loop interaction, this strategy allows rapid adaptation to new surgical contexts while avoiding overfitting and costly retraining.
@@
We evaluate the proposed system in video-guided skull base surgery scenarios and benchmark its 3D tracking performance against a commercially available optical tracking system. Across three experimental trials, the hybrid vision-based method achieved a mean absolute tool-tip position error of $2.32\pm 1.10$ mm in the camera frame, with inter-frame yaw and pitch propagation discrepancies of $0.18\pm 0.25^{\circ}$ and $0.21\pm 0.30^{\circ}$, respectively. The system also supports hands-free anatomy registration through our proposed virtual-cursor interface, achieving sub-millimeter reprojection accuracy in most trials. Tool segmentation and anatomy registration can be completed within approximately two minutes, substantially reducing setup complexity relative to conventional tracking workflows. These results show that a modular, speech-guided embodied agent can provide accurate video-based spatial guidance while improving workflow integration for surgical navigation.
@@ II Related Work
Geometry-aware pose estimation: Estimating the pose of surgical instruments from video is challenging due to limited texture, specular reflections, occlusion, and near-symmetry [21, 12]. In the broader 6D pose estimation literature, model-based methods such as ZebraPose [20] recover pose by predicting dense correspondences between image pixels and a 3D model, followed by PnP. Recent approaches such as FoundationPose and Any6D aim to generalize to unseen objects, but typically rely on RGB-D observations or strong geometric priors that are difficult to obtain in highly magnified surgical video.
In surgical contexts, prior work has explored incorporating additional structure to improve robustness. For example, robot-assisted settings can leverage kinematic priors or temporal filtering [6], while image-based approaches such as Hasan et al. [7] estimate pose from segmentation-derived geometric primitives under simplified assumptions (e.g., cylindrical tool models). While effective, such assumptions may not generalize well to tools with varying profiles, such as surgical drills. In this work, we adopt a geometry-aware formulation tailored to monocular surgical videos. We combine (i) segmentation-derived 2D geometry, (ii) monocular depth cues anchored to registered anatomy, and (iii) a CAD-based model representation to estimate tool pose.
Figure 1: System overview of the embodied surgical agent. The surgeon interacts through a hands-free interface (speech-to-text) that issues high-level commands to the front end, which orchestrates live video streaming, tool/anatomy segmentation, pose tracking, and anatomy registration. Intermediate outputs (masks, pose hypotheses, and registered anatomy overlays) are persisted in a streaming memory and can be retrieved on demand to support iterative refinement and rapid task switching without disrupting the surgical workflow. The back end executes modular perception and geometry components—including promptable segmentation, temporal mask propagation, surgical navigation, and pose/registration solvers—to produce stable tool state interpretation and anatomy-aligned navigation overlays in the endoscopic view.
@@ III Methodology
In this section, we describe the system architecture and interactive workflow of the proposed embodied surgical agent framework. The system is designed to support natural human–AI collaboration, where the surgeon surgeon interacts with the system through verbal commands, while the agent interprets intent and triggers specialist vision modules to perform the requested operation by invoking specialist vision modules directly on the live video stream.
III-A Overview
The framework operates as an interactive system that integrates language reasoning and visual perception (Fig. 1). A lavalier microphone captures real-time speech input from the surgeon, which is transcribed using Whisper speech recognition model. The transcribed query is processed by an agent, which performs stepwise reasoning to interpret the surgeon’s intent and generate structured action commands. These actions are executed by a set of specialist vision models. The system provides audio feedback through text-to-speech synthesis model, thereby enabling continuous bidirectional interaction. This design enables a conversational interface in which the surgeon can dynamically request and refine visual tasks on live endoscopic video streams. The interaction begins with the segmentation and labeling of the active surgical instrument followed by automated tracking of tip of the surgical instrument. Subsequently, the instrument tip acts as a persistent embodied spatial interaction pointer within the scene. This tracked virtual pointer enables downstream modules including anatomy segmentation, anatomy registration, vision-based surgical tool pose tracking, and anatomy-aware surgical navigation with critical-structure overlays. The following subsections describe the design of each component.
III-B Surgical Tool Segmentation
We adopted the speech-guided tool segmentation pipeline [1] for human-in-the-loop, segmentation and labeling of the active surgical instrument. This pipelines integrates text-promptable GSAM [18] model for generation of mask proposals for instrument and CUTIE [4] for video object tracking. However, a major practical limitation of this pipeline was that the requirement of instrument to remain stationary while the surgeon selected the desired mask, since tracking could begin only after instrument segmentation.
To address this limitation, we introduce a streaming memory module that buffers incoming frames while the surgeon interacts. Once the segmentation masks are selected for the initial frame $t_{0}$, the module performs sample a small set of frames uniformly up to the current frame $t_{n}$ and used them go seed a catch-up propagation step. This synchronizes the tracker with the live scene and allows real-time mask propagation to resume without requiring the tool to remain stationary during selection.
III-C Tip Point Tracking
@@
Once the instrument is tracked spatially, we utilize the segmentation mask to derive the tip of the instrument which serves as a dynamic spatial anchor to enable subsequent interaction with the surgical scene. Let $m_{t}$ denote the binary tool mask at time $t$, and let $\mathbf{e}_{t}^{(1)}$ and $\mathbf{e}_{t}^{(2)}$ denote the two extrema of the mask projection along its first principal axis. At initialization, the endpoint closer to the image boundary is treated as the tool shaft and the opposite endpoint as the tip. For subsequent frames, the tip is selected as the endpoint closest to the previous estimate:
@@
This suppresses axis flips and preserves directional consistency over time. To improve robustness under partial visibility of the tool base, we further apply a boundary-cropping step. We iteratively remove parts of the mask on the base side and recompute PCA until the base is no longer boundary-truncated and locate within the tool mask. This yields a more stable principal-axis estimate under partial views.
III-D Interactive Anatomy Segmentation
Figure 3: Interactive anatomy segmentation and refinement with event-triggered prompt retrieval. A voice command initiates anatomy segmentation while the system buffers the drill-tip trajectory. When the surgeon says “Done,” the buffered trajectory is converted into spatial prompts for SAM to generate an initial mask. Refinement is then triggered by a “Click” command on regions to remove, which adds prompts for mask update. The final anatomy mask seeds CUTIE for temporal propagation. Blue arrows indicate segmentation-mode prompting and trajectory retrieval, while purple arrows indicate refinement interactions and SAM updates.
We use the tracked tool tip as a dynamic interactive virtual cursor to segment anatomical regions within the surgical scene. Previous work on interactive segmentation focused on using depth estimates to simulate a virtual click [1], which are then used as point prompts for interactive VFMs, such as SAM [9], to segment an anatomical region. This approach relies on monocular depth estimates from DepthAnything [26], which we found to be temporally sensitive (see Fig. 13) for simulating virtual virtual clicks, which can lead to over- or under-segmentation.
To address this limitation, we introduces a region-constrained approach, which encodes user intent in image space to segment anatomical regions. As the user hovers the instrument over a region of interest, we collect a trajectory $\mathcal{T}=\{\mathbf{u}_{i}\}_{i=1}^{N}$, where $\mathbf{u}_{i}\in\mathbb{R}^{2}$ denotes the projected tool tip location in the image plane. A bounding box $B=[x_{\min},x_{\max}]\times[y_{\min},y_{\max}]$ is computed from $\mathcal{T}$, defining a constrained interaction region. We then sample a set of point prompts
@@ III-F Surgical Navigation
The interactive module enables the surgeon to dynamically select, update, and toggle anatomical structures during the procedure. Combined with speech-driven control, this supports continuous, hands-free interaction, allowing the visualization to adapt in real time to the surgical context and task requirements.
III-G Spatial Pose Tracking
We formulate pose estimation as a constrained geometric tracking problem. Rather than relying directly on image appearance, we combine multiple complementary cues, including segmentation masks, depth estimation, and geometric priors from a CAD model, to obtain a stable and physically consistent pose estimate. We first introduce a depth-based formulation and then refine it with additional 2D constraints to improve robustness and temporal consistency.
Mesh Geometry Primitive Extraction
[…]
These primitives provide a canonical geometric representation of the tool.
Depth-Derived Pose Tracking
The pose of an elongated and symmetric tool can be represented using a skeleton model defined by its tip position and principal axis. The depth-derived pose tracking approach consists of: (1) obtaining reliable depth estimation, (2) back-projecting the 2D tip position into 3D, and (3) computing the 3D principal axis via PCA on the back-projected tool mask.
Given a tool mesh and a predefined tip location on the mesh, these components allow us to estimate the tool pose in the camera frame.
Fore-ground Conditioned Depth Retrieval: Accurate 3D reconstruction of the tool axis and tip requires reliable foreground depth. We found stereo-based methods, including IGEV++ [25] and FoundationStereo [23], to degrade under high magnification because of the narrow effective baseline. In contrast, the fore-ground conditioned monocular depth model provides denser and more stable predictions, but only up to relative scale (Figure 6). To reliably recover metric scale for pose tracking, we fuse the monocular depth estimate from DepthAnything-v2 (DA2) [26] with registered anatomy depth from the visible anatomical regions, $S_{f}^{\mathrm{reg}}(\mathbf{u})$.
Figure 6: Stereo vs. fore-ground conditioned monocular depth for surgical tool reconstruction. Stereo depth estimates degrade under narrow baseline, leading to noisy 3D reconstructions, whereas DepthAnything v2 constrained to the segmented fore-ground produces denser, more consistent depth and a smoother tool point cloud.
Let frames be indexed by $t\in\{1,\dots,N\}$, and pixels by $\mathbf{u}=(u,v)$. Foreground mask is defined as the union of all propagated object masks:
$M_{f}^{\mathrm{fg}}(\mathbf{u})=\bigvee_{k\in\mathcal{O}}M_{f}^{(k)}(\mathbf{u}),$
where $\mathcal{O}$ denotes the set of tracked objects, and $M_{f}^{(k)}$ is the mask corresponding to object $k$.